OpenStudy (anonymous):

Need help, can someone verify my calculations for nuclear fusion equation? question inside

OpenStudy (anonymous):

Mass of deuterium = 2.014102u Mass of tritium = 3.014049u Mass of helium = 4.002603u Mass of neutron = 1.008664u Mass defect = (2.014102u + 3.014049u)-(4.002603u + 1.008664u) = 0.018884u =0.018884 x 931.5 MeV =17.590446 MeV per nuclei =17.590446x10^6 x 1.602x10^-19 =2.817989449x10^-12 joules per nuclei. Therefore in a 50-50 mix of deuterium and tritium where 5kg of tritium is used, there would be. =Number of moles(tritium)=mass/molar mass = 5/(3.014049 x 1.6606x10^-27 =10x10^26 moles of tritium Therefore 10x10^26 moles of deuterium are also used =Mass(deuterium)=number of moles x molar mass = 10x10^26 moles x (2.014102 x 1.6606x10^-27) =3.3412 kg of deuterium. So for 8.3412kgs worth of fuel we can achieve Number of particles = number of moles x avogardo’s number = 10x10^26 x 6.022x10^22 = 6.022x10^49 nuclei = 17.590446 MeV x 6.022x10^49 = 1.059296658x10^51 MeV =1.059296658x10^51 MeV x 1.602x10^-19 = 1.696993246x10^32 Joules of energy released.

OpenStudy (anonymous):

I believe i may have done something wrong, but i cant find it. My example for my assignment seems to have created something with more energy than the sun.

OpenStudy (anonymous):

the only thing that 's wrong is the lines: "=Number of moles(tritium)=mass/molar mass = 5/(3.014049 x 1.6606x10^-27" 5 should be replaced by 5000 (grams). This will translate into increasing the final answer by a magnitude of 3.